Calories protein( in grams) Vitamin C (in milligrams)

Food A

40 10 20
Food B 500 5 50
Food C 200 2 300

Three foods have the nutritional content per ounce shown in the table to the right. If a meal consisting of the three foods allows exactly 740 ?calories, 17 grams of? protein, and 930 milligrams of vitamin? C, how many ounces of each kind of food should be? used?
Answer #1

Let x, y, z ounces of Food A,B,C respectively be used. Then 40x+500y+200z = 740…(1), 10x+5y+2z = 17…(2) and 20x+50y+300z = 930…(3).

The augmented matrix of this linear system is A (say) =

40

500

200

740

10

5

2

17

20

50

300

930

To solve the above linear system, we will reduce A to its RREF as under:

Multiply the 1st row by 1/40

Add -10 times the 1st row to the 2nd row

Add -20 times the 1st row to the 3rd row

Multiply the 2nd row by -1/120

Add 200 times the 2nd row to the 3rd row

Multiply the 3rd row by 1/280   

Add -2/5 times the 3rd row to the 2nd row

Add -5 times the 3rd row to the 1st row

Add -25/2 times the 2nd row to the 1st row

Then the RREF of A is

1

0

0

1

0

1

0

1/5

0

0

1

3

It implies that x = 1, y = 1/5 and z = 3. Thus, 1, 1/5 and 3 ounces of Food A,B,C respectively should be used.